OpenStudy (anonymous):

Joey’s sock drawer is unorganized and contains 7 black dress socks, 7 black ankle socks, 6 brown dress socks, and 2 brown ankle socks. What is the probability that Joey will blindly reach into his sock drawer and pull out a sock that is brown or a dress sock? 4 d. A. 15/22 B. 4 C. 3/11 D. 21/22

OpenStudy (cwrw238):

probability of brown sock = 8/22 probability of dress sock = 13/22 can u continue from here?

OpenStudy (cwrw238):

with combinations of probabilities , ' or ' = + , and = '*'

OpenStudy (anonymous):

I got it. Thanks.